If you are finding the line perpendicular to x + 3y = 5, what is the slope of the line?
Answer by stanbon(75887)

If you are finding the line perpendicular to x + 3y = 5, what is the slope of the line?
==========
Slope of the given line:
3y = -x+5
y = (-1/3)x + (5/3)
----
slope = -1/3
----
Slope of a perpendicular line: 3
======================================
Cheers,
Stan H.